OpenStudy (rocklionmba):

Diana received 65 points on a project for school. She can make changes and receive three-tenths of the missing points back. She can do this 10 times . Create the formula for the sum of this geometric series, and explain your steps in solving for the maximum grade Diana can receive.

OpenStudy (rocklionmba):

@dumbcow @aaronq

OpenStudy (dumbcow):

\[35(\frac{3}{10} + \frac{3}{10}^2 +...\frac{3}{10}^{10})\]

OpenStudy (rocklionmba):

this is diverging... am I right?

OpenStudy (dumbcow):

no converging because 3/10 raised to big number goes to zero

OpenStudy (dumbcow):

formula for sum: \[\frac{1-r^n}{1-r}\] in this case r = 3/10

OpenStudy (rocklionmba):

and then n would be 10 right?

OpenStudy (dumbcow):

yes

OpenStudy (rocklionmba):

(btw I leave and come back of a glitch in openstudy that does that) so then we first have to the r^n

OpenStudy (rocklionmba):

hoe do we get that? im bad with fractions

OpenStudy (dumbcow):

use a calculator .... (3/10)^10

OpenStudy (rocklionmba):

0.0000059049

OpenStudy (rocklionmba):

I think I got it from here. thanks

OpenStudy (dumbcow):

ok :)